OpenStudy (anonymous):

Let \(a_{n} = \frac{1}{n^{2}}\) then \( r = \lim_{n \to \infty} |\frac{a_{n}}{a_{n+1}}|= \lim_{n \to \infty}|\frac{(n+1)^{2}}{n^{2}}|\) Here why we behave to \(\frac{1}{n^{2}}\) like as its a \(\frac{n^{2}}{n^{2}}\) ?

OpenStudy (anonymous):

in second lim... why we have \[(n+1)^{2}\] ?

OpenStudy (anonymous):

i think it should be \[1\]

OpenStudy (anonymous):

actually i know when we see a n we should do it n+1 but where is our 1 from \[\frac{1}{n^{2}}\]

OpenStudy (kinggeorge):

You wrote: \[r = \lim_{n \to \infty} \left|\;\frac{a_{n}}{a_{n+1}}\right|\]It should be\[r = \lim_{n \to \infty} \left|\;\frac{a_{n+1}}{a_{n}}\right|\]

OpenStudy (anonymous):

hmm in original question its like i wrote, maybe our mentor write it wrong... but my question is how we can use quotient rule to \[\frac{1}{n^{2}}\]

OpenStudy (anonymous):

they get above \[\frac{(n+1)^{2}}{n^{2}}\] but how its possible ? :(

OpenStudy (kinggeorge):

I've got leave after this, but if you use \[r = \lim_{n \to \infty} \left|\;\frac{a_{n+1}}{a_{n}}\right|\]you can get to the following (after some simplification).\[r=\lim_{n \to \infty} \left|\;\frac{n^2}{(n+1)^2}\right|=\lim_{n \to \infty} \;\frac{n^2}{(n+1)^2}\]Use l'hopitals rule on this, and you get a solution for \(r\).

OpenStudy (anonymous):

ok thanks George :)